A metallurgist needs to produce 10kg of an alloy which is 65% copper and 35% tin. He has an alloy which is 50% copper and 50% tin. How Much of this 50-50 alloy must he melt together with pure copper to make the desired alloy? How much copper must he use? If necessary, round each answer to the nearest tenth of a a kilogram.
Answer by mananth(16946)

Copper
percent ---------------- quantity
Metal alloy I 50 ---------------- x Kg
Pure copper 100 ------ 10 - x Kg
Mixture 65 ---------------- 10
600
50x+100 (10-x) =65*10
5000x+10000(10 -x)=65000
5000x+ 100000 -10000x = 65000
5000x - 10000 x = 65000 - -100000
-5000 x = -35000
/ -5000
x = 7 Kg 50.00% Metal alloy I
3 Kg 100.00% Pure copper

m.ananth@hotmail.ca